[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Fri Jul 21, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Fri Jul 21 19:59:55 UTC 2017


HSU Shift Log for Fri Jul 21 2017 15:59:49 GMT-0400 (EDT)

Shift: H

Forecaster: E Blake

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues: The AMSU size estimates in the SWHG directory are  
missing. (nothing in the fields last 24 hours)

Miscellaneous issues: GFS initialization at 12Z for 9-E looked very bad.  
The pressure fields was all messed up. Something may have gone wrong in the  
DA system. Might be worth mentioning to EMC. It also affected HWRF/HMON too.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170721/dc38317d/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list